OpenStudy (kymber):

The angle of elevation of a ladder leaning against a wall is 72°. If the ladder is 20 feet long, how far is the base of the ladder from the wall? I got 19.02, is that right? (Pleeeeease be right ಠ෴ಠ)

OpenStudy (anonymous):

cos72 = x/20 20cos72 = x 20(0.309) = x x = 6.18 ft

OpenStudy (anonymous):

\[20\text{ Cos}[72{}^{\circ}]=5 \left(-1+\sqrt{5}\right)=6.2 \text{ feet}\]

OpenStudy (kymber):

Waha. Thaanks